IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

  • Out of the office.

Serge Chastel

  • Moved ippMonitor to ippdb03 (ippMonitor.ipp.ifa.hawaii.edu is its alias)
  • Fixed warpProcessed_Images (Note: new configuration is not changed when deploying new ipp version)
  • Limited some requests from ippMonitor when there is no filtering
  • Spent a lot of time with psphot and flipping images for zoology: discovered that psphot had a random part once all experiments have been run

Heather Flewelling

  • false detections
  • ifaps1 stress testing rebuilt raid
  • czar on 3 days
    • investigated why burntool has problems...
  • liason
    • investigating TD's stacks (why not deep enough/not enough area, need more info)
    • investigated RW's missing exps
    • investigated SAS and LAP dvodb for Roy,
  • addstar - rsync complete - need to restart addstar on ipp004/ipp005.

Roy Henderson

  • three day week (vacation on Thurs/Fri)
  • ippToPsps development:
    • now only importing relevant columns from smf files into the database. Speeds things up substantially
    • changed logic so that we can abort a batch before importing the smf if there are no IDs to be found in DVO
    • improved metrics program for monitoring loading, including more robust statistics as well as a count of failed as well as pending items
    • now correctly logging to the database if loading to the datastore fails
  • other
    • czarpoll now querying replicated gpc1 database on ippdb03, rather than production master on ippdb01
    • updated the PSPS news page, ready for new SAS data and imminent 3PI data
    • continued to monitor loading of LAP data

Mark Huber

  • MD02.V3 refstack completed, verified and distributed. MD10, 01, 02 evaluation continues.
  • Increased difference image detections in some exposures look to be caused by glitch from tracking or aberration. Sample set selected and updated to visually inspect images.
  • simtest IPP->PSPS test configuration debugging continued.

Bill Sweeney

  • Wrote new perl function to implement safe stdio redirection. Modified many of the ippScripts to use it.
  • 1 day as official processing czar and many hours "assisting". Since the number of nodes with free disk space is currently small the IPP has more errors to repair.
  • Postage stamp server monitoring and repair: We had quite a bit of activity last week. The error rate was much higher than expected so I investigated. The problem was primarily caused by ipp005 outage from several weeks ago. Components were marked as gone when they should not have been. Repaired these.
  • At the end of the week it was decided that the update processing was placing too large a load on the database and processing nodes. Tuned some parameters for both of those.
  • Made postage stamps from some special processing that Gene did for Niall.
  • Started looking at the 450% worse false streak reporting that is affecting M31 observations this year. The M31 team considers the data unusable as is. It does not seem to be due to the new template or tessellation. This Summer's counts were worse than 2010 before we switched.

Chris Waters

  • Nightly Science: extended cleanup timeout period to prevent the cleanup command from not fully completing.
  • Diskspace: Finished coding and testing of first instance retargetting shuffle. This should free up space on wave2/wave3 nodes, as well as allowing processes to run somewhat faster due to decreased I/O. Need to increase the speed, as it appears to not be running quite fast enough. This is also finding inconsistent nebulous objects, and I have a script in place to repair these.
  • LAP: continued LAP processing, with some minimal planning for the next revision.